OpenStudy (anonymous):

1. Write the equation of the line with a slope of 3/7 passing through the point (9, -2) 2. Write the equation of the line parallel to X = -16 and passes through the point (9, -10) 3. Calculate the equation of the line between the points (-6, -1) and (2, -1) 4. Determine if the lines 9x -2y = 16 and 2x -9y = -18 are parallel, perpindecular, or neither

OpenStudy (hba):

1) \[y_2-y_1=m(x-x_1)\] Where m=3/7 \[(x_1,y_1)=(9,-2) \]

OpenStudy (hba):

Please post a single question in a single post. Thanks :)